OpenStudy (anonymous):

a steel cable is 45m long and has a radius of 2.3cm. how much force is required to stretch it 0.15mm? young's modules for steel is 210E9 N/m^2.

OpenStudy (phebe):

u still here

OpenStudy (sidsiddhartha):

do u know anything about young's modulus?

OpenStudy (sidsiddhartha):

Y = Stress/Strain ever seen this formula? @cro825

OpenStudy (shamim):

Y=FL/(Al) is ur formula needed

OpenStudy (anonymous):

and that translates to f=YAL/Lo right?

OpenStudy (shamim):

F=YAl/L

OpenStudy (anonymous):

thanks i will try that

OpenStudy (shamim):

u r most welcome

OpenStudy (shamim):

if u r pleaded on my solution then u can give me s medal

OpenStudy (anonymous):

how i do that?

OpenStudy (shamim):

just click on right side of my name shamim

OpenStudy (anonymous):

idk im getting 1.6E6 but not correct

OpenStudy (shamim):

2.3cm=?m

OpenStudy (shamim):

0.15mm=?m

OpenStudy (anonymous):

.0023

OpenStudy (shamim):

no

OpenStudy (anonymous):

.023

OpenStudy (shamim):

yes

OpenStudy (shamim):

now 0.15mm=?m

OpenStudy (anonymous):

.00015

OpenStudy (shamim):

right

OpenStudy (shamim):

now use those data into ur formula

OpenStudy (anonymous):

thought i was but i will again

OpenStudy (anonymous):

so 210E9(.023)(.00015/45) ?

OpenStudy (shamim):

A=cross sectional area of given wire=πr^2

OpenStudy (anonymous):

crap thought i could use the 2.3 cause it was not the dia i knew i was to find proper A and had the formua with me

OpenStudy (anonymous):

sodo i use the cm or meter conv. in the area?

OpenStudy (shamim):

u hv to use m^2 for area

OpenStudy (shamim):

A=πr^2=3.1415*0.023^2m^2

OpenStudy (anonymous):

ok i had that right then i over think at times so yes 210E9(1.66E-3)(.00015/45) = 1.2E3 for 2sf and was correct. thank you so much

OpenStudy (anonymous):

got time for a similar?

OpenStudy (shamim):

welcome

OpenStudy (shamim):

anyway which grade u r

OpenStudy (shamim):

just for my curiosity

OpenStudy (anonymous):

college. pretty slow arent i lol

OpenStudy (shamim):

no no

OpenStudy (anonymous):

but im 40 and just stared back a yr ago

OpenStudy (shamim):

its ok

OpenStudy (shamim):

hm

OpenStudy (shamim):

r u frm usa

OpenStudy (anonymous):

yes i am in georgia

OpenStudy (shamim):

so georgia is a state in usa. right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

still around

OpenStudy (anonymous):

tanks again the help me solve the next problem i had to b seeing were i went wrong on this one